OpenStudy (anonymous):

Last question of the night!!! A circle in the first quadrant with center on the curve y = 2x2 − 27 is tangent to the y-axis and the line 4x = 3y. The radius of the circle is m/n where m and n are relatively prime positive integers. Find m + n.

OpenStudy (anonymous):

y = 2x^2

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

The center is on the curve 2x^2? Not on the curve 2x^2-27?

OpenStudy (anonymous):

the center is on the curve y = 2x^2-27

OpenStudy (anonymous):

i forgot to put the "^" in the original question xD

OpenStudy (anonymous):

Ah. Here's my thought process... For a given point on that curve, the radius of the circle will be DEFINED by the perpendicular distance to the y axis. If that radius is the same as the nearest point on y = (4/3)x, then that's our circle.

OpenStudy (anonymous):

For a given point (x,y) on the line, the distance to the y axis is just x. The distance to the line y = (4/3)x is: |4x -3y|/sqrt(25) = |4x-3y|/5 From here:

OpenStudy (anonymous):

Setting those equal: x = |4x-3y|/5 Now, y is a function of x. y = 2x^2 -27, so substituting in gives: x = |4x -3(2x^2-27)|/5 =|-6x^2 +4x+81|/5

OpenStudy (anonymous):

Solve for x.

OpenStudy (dumbcow):

however, the circle defined above which satisfies the conditions is Not in the 1st quandrant, the center would lie in the 4th quadrant.

OpenStudy (anonymous):

Oh really?

OpenStudy (anonymous):

Well damn...

OpenStudy (dumbcow):

wait nevermind i didn't look at the neg case of absolute value it can also be in 1st quadrant wow this problem is giving me fits for some reason